OpenStudy (anonymous):

Given the point (4, 5) and the slope of 6, find y when x = 24. 118 144 125 149

OpenStudy (anonymous):

first find equation of line as y-5=6(x-4) solve to get y=f(x) y=6x-19 now put x=24 y=6(24)-19 y=125 => c is correct

OpenStudy (anonymous):

thank you @Aditya9826042721